find the probability that at least 2 juniors are picked for a 3 member board when 8 juniors and 5 seniors are available.
Answer by stanbon(75887)

find the probability that at least 2 juniors are picked for a 3 member board when 8 juniors and 5 seniors are available.
----------------------------------------
# of ways to succeed: 8C2*5C1 + 8C3 = 140+56 = 196
# of possible outcomes: 13C3 = 286
------
P(2<= x <=3) = 196/286 = 0.6853
=================
Cheers,
Stan H.